# Generative Point Cloud Registration

### v1

@@ 5 Conclusion
We have introduced a novel 3D registration paradigm, generative point cloud registration, which effectively leverages advanced 2D generative models to augment geometry-only 3D registration. To this end, we developed Match-ControlNet, a matching-specific variant of ControlNet designed to synthesize paired RGB images for both source and target point clouds. By integrating depth-conditioned generation from ControlNet, coupled conditional denoising, and coupled prompt guidance, these generated RGB image pairs preserve both 2D-3D geometric consistency and cross-view texture consistency, thereby facilitating high-quality 3D matching. Notably, our generative framework is general and can be incorporated into a variety of registration methods to improve their performance. Extensive experiments demonstrate the effectiveness of the proposed paradigm.
Impact Statement
This paper presents work whose goal is to advance research in point cloud registration. While potential societal implications may exist, we do not find any that necessitate explicit discussion here.
@@ Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
Appendix A More Quantitative Analysis
To further validate the effectiveness of our Generative Point Cloud Registration paradigm, we integrate the handcrafted geometric descriptor, FPFH, into our framework, forming Generative FPFH. As shown in Table 4, this generative variant achieves a significant performance improvement over the baseline FPFH on the 3DMatch benchmark dataset, regardless of whether DINOv2 encoding or Stable-Diffusion encoding is used.
Table 4: Comparison of the methods on rotation, translation, and Chamfer distance on 3DMatch (zeng20173dmatch) benchmark dataset.
Rotation (deg) Translation (cm) Chamfer (mm)
[…]
Methods 5 10 45 Mean Med. 5 10 25 Mean Med. 1 5 10 Mean Med.
FPFH (rusu2009fast) 69.1 82.9 91.2 15.0 3.1 25.8 53.9 75.1 37.4 9.1 52.5 74.2 79.2 57.6 0.9
[…]
Generative FPFHSD 88.2 94.7 96.9 6.1 2.0 37.6 68.9 85.5 18.5 6.6 66.9 85.7 88.8 41.2 0.6
Improvement $\uparrow$ 19.1 11.8 5.7 8.9 1.1 11.8 15.0 10.4 18.9 2.5 14.4 11.5 9.6 16.4 0.3
Appendix B More Visualization Results of Match-ControlNet
In Fig.7, we present additional visualization results of the RGB image pairs generated by our Match-ControlNet, along with the corresponding colorized source and target point clouds. Furthermore, Fig.8 visualizes the image pairs generated by our zero-shot Match-ControlNet, while Fig. 9 showcases image pairs produced by the finetuned Match-ControlNet.
Figure 7: More visualization results of the generated RGB image pairs and the formed color source and target point clouds.
Figure 8: Source and target image generation via zero-shot Match-ControlNet without any finetuning.
Figure 9: Source and target image generation via finetuned Match-ControlNet.

### v2

@@ 5 Conclusion
We have introduced a novel 3D registration paradigm, generative point cloud registration, which effectively leverages advanced 2D generative models to augment geometry-only 3D registration. To this end, we developed DepthMatch-ControlNet and LiDARMatch-ControlNet, two matching-specific variants of ControlNet designed to synthesize paired perspective-view/panoramic RGB images for both source and target point clouds. By integrating depth/range map-conditioned generation from ControlNet, coupled conditional denoising, and coupled prompt guidance, these generated RGB image pairs preserve both 2D-3D geometric consistency and cross-view texture consistency, thereby facilitating high-quality 3D matching. Notably, our generative framework is general and can be incorporated into a variety of registration methods to improve their registration performance. Extensive experiments on both depth camera-based/LiDAR-based 3D registration tasks demonstrate the effectiveness of the proposed framework.
@@ ACKNOWLEDGMENTS
This research is supported by MOE AcRF Tier 1 Grant of Singapore (RG12/22) and by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).